OpenStudy (anonymous):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 9x2 − 6x + 1 square feet, what is the length of one side of the room? (3x + 1) feet (9x − 1) feet (3x − 1) feet (9x + 1) feet

OpenStudy (goformit100):

(3x − 1) feet

OpenStudy (anonymous):

How did you get it?

OpenStudy (anonymous):

@goformit100 how did you get that

OpenStudy (rsadhvika):

area = l^2 = 9x2 − 6x + 1 = (3x)^2 - 2(1)(3x) + 1 = (3x - 1 )^2 l = sqrt ( (3x-1))^2 = 3x-1

OpenStudy (anonymous):

Thanks :D

OpenStudy (rsadhvika):

yw !(: